The following information is used for the rest of this page’s questions:
In order to buy a new warehouse, Phillips Corp. took out a $2,000,000, 12% annual rate mortgage from Burton Bank on January 1, 2018. Philips agreed to make monthly payments of $25,000 on the mortgage at the end of each month until the balance is eliminated.

How much interest is incurred on the mortgage in February 2018, the second month?

After the payment, what will the mortgage balance be on March 31, 2018, the end of the third month?



After six monthly payments, Phillips refinanced the mortgage by signing a new 15-year mortgage at a 6% annual rate. An additional $10,000 of bank fees were added to the principal balance in order to cover the refinancing costs. What equal payments must Phillips now make at the end of each month in order to satisfy the new mortgage? (Round to the nearest dollar)
Answer #1

Answer (1):

Jan 2018:

Interest = 2000000 * 12%/12 = $20,000

Repayment = 25000 - 20000 = $5,000

Ending balance = 2000000 - 5000 = $1995000

Feb 2018:

Interest = 1995000 * 12%/12 = $19950

Hence:

Interest incurred on the mortgage in February 2018, the second month = $19,950

Answer (2):

We calculate below amortization schedule for first 6 months:

A Month Beginning Balance Monthly Payment Repayment Ending Balance =C-B I = A-D Jan-18 Feb-18 Mar-18 Apr-18 May-18 Jun-18 $2,

Hence:

Mortgage balance be on March 31, 2018 = $1,984,850

Answer (3):

From answer 2 above, we find outstanding mortgage balance after six months = $1,969,240

With $10,000 of bank fees, total loan amount = $1,969,240 + 10000 = $1,979,240

Monthly interest = 6%/12 = 0.5%

Number of monthly installments = 15 * 12 = 180

Hence:

Monthly payments = PMT (rate, nper, pv, fv, type) = PMT (0.5%, 180, -1979240, 0, 0) = $16,701.95

Phillips now must make equal payments at the end of each month in order to satisfy the new mortgage = $16,701.95